OpenStudy (anonymous):

Thirteen plums weigh as much as two apples and one pear. Four plums and one apple have the same weight as one pear. How many plums have the same weight as one pear?

OpenStudy (anonymous):

I have no idea how to solve this

OpenStudy (anonymous):

13p = 2a + x 4p + a = x p = plums, a = apples, x = pears you need to find an equation with onle plums(p) and pears(x) by solving for a on the second equation: a = x - 4p and substituting the first equation: 13p = 2(x - 4p) + x 13p = 2x - 8p + x 13p = 3x - 8p 21p = 3x (21 plums = 3 pears, but the questions is plums to ONE pear, so solve for x) x = 7p 7 plums = 1 pear

OpenStudy (anonymous):

Wow!!! Thank you so much!!!!

OpenStudy (anonymous):

no problem